Reports of 58720
AdV-DET (Commissioning)
masserot - 16:20 Thursday 23 May 2024 (64333) Print this report
Comment to B1s_PD1_112MHz phi computation (64309)

The B1s_PD1_112MHz_mag threshold has been updated  to 0.005  the 2024-05-21-12h47m16-UTC .

Looking at the B1s_PD1_112MHz_{mag,phi} signal trend , trend plot and its zoom,  one can find

  • a  big jump on the B1s_PD1_112MHz_phi signal  when the B1s_PD1_112MHz mag reached the threshold 
  • but  far enough and short (see the raw plot )before enabling the  phase noise subtraction using the B1s_PD1_112MHz phi channel .

This issue has been understand : it s present only on the ACL_OP_CH keyword using the operator  "mag|phi_umwrap_n", not in the  simple "mag|phi" operator . This issue has been fixed in the Acl release v1r25p17

Images attached to this comment
AdV-COM (AdV commissioning (1st part) )
mantovani - 15:36 Thursday 23 May 2024 (64332) Print this report
Comment to Range decrease to 45Mpc because of etalon (64328)

I think it is hard to address all the lost MPc to etalon. I fear that an other drift (not clear of which parameters) is adding on the top.

As it is visible for NI te < 20.08, the incriminated drift, the behavior of BNS and DIFFp tx has changed behavior and it can point to an other monotonic drift adding on the top of the Etalon drift.

In any case improving the stability of Etalon will better disantangle the other drifts of the machine helping the understanding

Images attached to this comment
Virgo Runs (O4b)
gherardini - 14:59 Thursday 23 May 2024 (64331) Print this report
Operator Report - Morning shift

The ITF unlocked from science at 7:29UTC most probably because of a failure of the SNEB vertical control (see plot), relocked after the recovering of SNEB controls, science mode started at 8:34UTC.

 

Sub-system reports

SBE
7:29UTC the SNEB vertical control opened by the guardian, controls properly closed after recover the vertical position with F0 step motor.

Images attached to this report
Virgo Runs (O4b)
menzione - 6:03 Thursday 23 May 2024 (64330) Print this report
Operator Report - Night shift

At 21:00 UTC I found ITF just exit from Science mode due to the unlock of the SQZ. Properly relocked. Back in Science at 21:03 UTC.
It remained in Science Mode for the whole shift and the Horizon increased up to it's standard range. 

Guard tour (UTC):
21:06 - 21:34
23:05 - 23:36
01:31 - 02:00
03:30 - 04:00

Images attached to this report
Virgo Runs (O4b)
Sposito - 23:04 Wednesday 22 May 2024 (64329) Print this report
Operator Report - Afternoon shift

I found ITF in SCIENCE; 21.00 UTC SQZ disengaged. 21:03 UTC ITF back in SCIENCE.

Guard Tour (UTC): 
17:14 - 17:40
19:11 - 19:41

Images attached to this report
AdV-COM (AdV commissioning (1st part) )
mwas - 21:16 Wednesday 22 May 2024 (64328) Print this report
Range decrease to 45Mpc because of etalon

Figure 1. After the day mostly unlocked yesterday there was a large transient on etalon temperature control. This resulted in a decrease in optical gain, decrease in BNS range, and the DIFFp TX error signal based on the DARM line moving away from zero. It underlines the importance of implementing the scheme tested at the end of O3 to keep the etalon control stable during unlocks.

Images attached to this report
Comments to this report:
mantovani - 15:36 Thursday 23 May 2024 (64332) Print this report

I think it is hard to address all the lost MPc to etalon. I fear that an other drift (not clear of which parameters) is adding on the top.

As it is visible for NI te < 20.08, the incriminated drift, the behavior of BNS and DIFFp tx has changed behavior and it can point to an other monotonic drift adding on the top of the Etalon drift.

In any case improving the stability of Etalon will better disantangle the other drifts of the machine helping the understanding

Images attached to this comment
Environmental Monitoring (Environmental Monitoring)
fiori, tringali, spinicelli, paoletti, napolano, derossi - 18:31 Wednesday 22 May 2024 (64326) Print this report
Music perfomance test

Yesterday between 13:46 and 13:59 UTC we performed a test of the music performance planned for this Friday night, during the EGO-Virgo open day. Loudspekers (two THUMP 15 Cassa Acustica a 2 vie with 15” woofer, 1000 Watt) are located as indicated in Figure 1, pointing West. Tracks of the real performance were played, same volume.

The time of the test was not the best since there was intense wind (30 to 40 km/h) and the ITF could not lock. The test was done anyway with the intention to look and environmental momitors. No evidence of excess noise was evident in the environmental probes (microphones and seismometers) in CEB and MCB. However, an excess of acoustic noise from wind gusts was noticeable in the CEB microphone up to about 200Hz (Figures 2,3,4) which might have hidden some music noise. Note that the excess acoustic noise peaking around 8AM UTC noticeable in Figure 3 was due to grass cutting.  The effect from the same wind gusts in the MCB mike is noticeable only up to 10 Hz or so, because of the larger infrastructure noise.

At 13:52:01 UTC the ITF unlocked from state 74 (CARM_MC_IR) and also the IMC went down (last Figure). This is a state with green to IR transition that is know to be sensitive to wind.

It has been decided to repeat the test anyhow in a commissioning slot tomorrow 17:30 LT, weather forecast foresee low wind and possibly ITF could be locked.

Images attached to this report
AdV-DAQ (Data collection)
masserot, verkindt - 18:17 Wednesday 22 May 2024 (64325) Print this report
Comment to Hrec latency : PyDAQ data collection updated (64247)

The attached plots show that

  • thank the migration of the SQZ_FLT frame collection on the SqzMoni server , there is less glitches on the FbmAlp server  (purple and green rectangle)
  • but few remains (red circle) without any time correlation with the PyUGF or SQZ_FLT latency glitches

All the latency glitches on FbmAlp server, related to the 2024-05-21-10h to 2025-05-22-10h time period have been analyzed

  • for the 2024-05-21 10h to 14h time period, several lock acquisition sequences : latency glitches due to FbsDet and ITF_LOCK servers
  • around 2024-05-21-15h30 : latency glicthes due the 2 FbsVac restarts
    • 2024-05-21-15h18m35-UTC    info sentenac     process:FbsVac started
    • 2024-05-21-15h28m01-UTC    info sentenac     process:FbsVac started
  • around 2024-05-21-15h40: latency glicthes due t0 the ITF_LOCK server
  • for the  2024-05-21  16h to 16h40 time period: latency glitches due the FbsDet (red rectangle), FbsAlp (orange rectangle) and SUS_PR node (yellow rectangle)
  • around 2024-05-21-18h24: latency glicthes due to SUS nodes
  • for the 2024-05-21 04h00 to 05h00 time period: latency glitches due the  SUS nodes (red rectangle), and one from FbmMain (green rectangle)

For  this time period analysis, over all these events only one is due to the FbmMain latency glitch , so as primary conclusion using FbmMain as frame provider for Hrec should help to reduce the number of latency glitches

Images attached to this comment
Infrastructures (Air Conditioning)
soldani - 15:56 Wednesday 22 May 2024 (64327) Print this report
HVAC tuesday maintenance_21.05.2024
Yesterday, during Tuesday maintenance,:
- carried out weekly checks for boiler, AHU and chillers in TB, CEB and MC
- refill of the LPG tank in Central area from 9.30 to 9.45 (UTC)
- replacement of the drainage pump in right-corner in Loc. AHU CEB
No checks or maintenance with interferences on hall temperature
Virgo Runs (O4b)
berni - 15:00 Wednesday 22 May 2024 (64324) Print this report
Operator Report - Morning shift

ITF found in Science mode.

At around 8:10 UTC there was a drop in Hrec_Range_BNS; then it started to fluctuate below the standar value (see attached plot). At 9:35 UTC there was another strong drop; in this occasion I removed Science mode for about two minutes.

The time of the drops seems to match the time of two local earthquakes:

 

Software

from 6:22 UTC to 6:52 UTC ITFOnCallMoni and DMSserver to take into account new configurations requested by experts.

 

Environment

Wind activity increasing during the shift.

 

Images attached to this report
Virgo Runs (O4b)
menzione - 6:49 Wednesday 22 May 2024 (64323) Print this report
Operator Report - Night shift

ITF in Science Mode for the whole shift.

Guard tour (UTC)
21:02 - 21:29
23:00 - 23:30
01:05 - 01:34
03:26 - 03:55

Images attached to this report
Virgo Runs (O4b)
Sposito - 23:00 Tuesday 21 May 2024 (64317) Print this report
Operator Report - Afternoon shift

I found ITF unlocked due to Bad Wheather

13:46-13:59 UTC V. Napolano music concert test

15:03 UTC ITF mode in COMMISSIONING. 15:07-15:31 UTC I started the OMC scan in CARM_NULL_1F as requested by R. Gouaty. 16:13-18:31 UTC ITF in LOW_NOISE_2. 

18:57 UTC I changed the etalon set point NI - 20.06 and WI 19.85 as requested by M. Mantovani.

19:02 UTC ITF back in SCIENCE

Guard Tour (UTC): 
17:02 - 17:29
19:01 - 19:29

No DMS events

Images attached to this report
AdV-COM (1/√f noise)
mwas - 21:21 Tuesday 21 May 2024 (64322) Print this report
Comment to CP thermal noise question (63854)

An example of excess thermal noise when mounting a mirror in a metal frame has been abserved at the AEI 10m prototype facility: https://dcc.ligo.org/LIGO-G2201622 .

The inferred substrate Brownian noise when the mirror is glued inside a metal frame 7e-18 m/rtHz. When the mirror is just clamped with screws the noise is even higher, and shows mechanical resonances (at a few kHz).

It is a very different case for the CP, but shows that a mount can change the substrate thermal noise to levels that would be sufficient to explain the mystery 1/f^{2/3} noise.

AdV-DET (Commissioning)
gouaty - 18:57 Tuesday 21 May 2024 (64320) Print this report
Fine adjustment of B2_QD2 Vertical galvo correction

Today, during the lock in CARM_NULL_1F around 16h00 utc, we performed some little adjustments with the picomotor on SIB2 folding mirror to bring the correction applied on the B2 QD2 vertical galvo closer to 0. The goal was to make the peak-to-peak fluctuations of the galvo corrections more symmetrical around 0 during the lock acquisition. This should reduce the risk of having the B2 QD2 galvo loop open during lock acquisition.

Images attached to this report
Optical characterization (Optical characterization)
sposito, bersanetti, mantovani, gouaty - 18:49 Tuesday 21 May 2024 (64319) Print this report
OMC scan in CARM NULL

Figure 1 shows the scan of the OMC performed today from 15h08 utc to 15h30 utc, after waiting for 1 hour in CARM_NULL_1F.

Figure 2 shows a zoom of this scan around a Free Spectral Range when the temperature was increasing.

In this FSR, assuming a calibration factor of about 4800 for B1_PD3, we have: ~50 mW on the first order mode, ~40 mW on the second order mode, ~130 mW on the third order mode, ~140 mW on the fourth order mode, around ~40 mW on each of the 5th and 6th order modes, ... At first sight it seems that most of the high order modes have comparable powers compared to the measurement performed one week ago ( https://logbook.virgo-gw.eu/virgo/?r=64257 ).

Focusing on the 2nd order mode which appears 4 times during the scan, its power is ranging between 33 and 43 mW which is quite comparable to the range of values found last week (although maybe a little bit higher in average, but I am not sure it is relevant given the fluctuations).

Images attached to this report
Environmental Monitoring (Environmental Monitoring)
fiori, ruggi - 18:11 Tuesday 21 May 2024 (64318) Print this report
Comment to Human noise injection test (64283)

As Paolo pointed out,  the PR drift that caused the unlock during the human noise injection test was most likely uncorrelated with the visitors presence, being instead one of the infamous "PR trips" affecting the ITF in those days (Figure 1) https://logbook.virgo-gw.eu/virgo/?r=63961 and  https://logbook.virgo-gw.eu/virgo/?r=63994.

On the other hand, the presence of people inside the CEB caused, as reported, noticeable drifts of suspensions (in particular IB suspension drifted by more than 20 micrometers) which cannot be tollerated. As reported, the presence of groups of people outside but in proximity of CEB causes an increased seismic noise in the approx. 1-5Hz band, which is presently suspected to correlate with excess noise in B1: https://logbook.virgo-gw.eu/virgo/?r=64287.

The conclusion is that, at least in the present ITF condition which is paricularly sensitive to anthropic seismic noise, the presence of visitors groups close to CEB while in science data taking shall be avoided.

Images attached to this comment
AdV-ISC (Automatic Alignment)
mantovani - 16:56 Tuesday 21 May 2024 (64316) Print this report
Etalon wp

The new point chosen, looking at the BNS and DiffP wp figure 1 and 2, are NI = 20.06 and WI =19.85.

Images attached to this report
Virgo Runs (O4b)
rolland - 15:27 Tuesday 21 May 2024 (64315) Print this report
Comment to Operator Report - Afternoon shift (64273)

After discussion, we agreed that the measurement is still meaningful to transfer the calibration from the PCal to the NI,WI mirror actuator in HP mode (then used to calibrate the BS +PR,SR mirror actuators). Doing the calibration in the new LOW_NOISE_1 mode looks ok (with the warning that calibration analysis should use the CAL_NI,WI_MIR_Z_CORR channels and not the Sc_NI,WI_MIR_Z_CORR channels).

Now, the CALI state DF_LN1calib_to_LN2ready is mainly doing nothing, instead of opening the NI,WI relays as it was done until now, since the NI,WI mirrors are now still used for the ITF control at this step. To be checked during the next calibration data taking that ITF does not unlock anymore after this measurement.

    def main(self):        ezca.put_index(self.index)        #log.info('Setting the NI,WI actuators to their standard state for LOW_NOISE ITF locks')        #self.message = "WI and NI coil relays set to 10 kOhm series resistor"        self.message = "WI and NI coil relays kept in HP mode"        #sb_10K()        log.info(self.message)        self.timer['SAT_RELAY'] = 1 #ramp of 5 s
AdV-DET (Commissioning)
masserot - 15:10 Tuesday 21 May 2024 (64314) Print this report
Comment to B1s_PD1_112MHz phi computation (64309)

A mistake was done for the B1s_PD1_112MHz phase computation , it was put before the rotation with 2xF56MHz LNFS phase instead of after .

This has been fixed at 2024-05-21-12h46m29-UTC by upgrading the SDB2_Photodiodes configuration and the restarting of the server

Virgo Runs (O4b)
berni - 15:00 Tuesday 21 May 2024 (64307) Print this report
Operator Report - Morning shift

ITF found in Science mode.

At 6:00 UTC I set MAINTENANCE mode; at 6:15 UTC I manually unlocked to perform the periodic task:

  • 6:05 UTC checks and refill TCS chiller and as requested by Ilaria I switched on the PIEZO-DRIVER;
  • 6:22 UTC OMC lock and scan in NI single bounce;
  • 7:10 UTC check of the IMC working point position;
  • 7:20 UTC TCS power checks;
  • 7:42 UTC TCS Therma Camera Reference;
  • 10:02 UTC FSR scan; the scan was interrupted by an unlock at the second iteration.

Other activities communicated in control room:

At 10:00 F. Nenci went in CB to verifty that all the operations were concluded and to switch off the light and close the main door.

At 10:20 UTC all the maintenance activities were concluded thus I set PREPARE_SCIENCE mode and I started to relock; after a few failed attempts at 11:52 UTC  I set BAD_WEATHER mode as the WIND activity was high.

At 12:47 UTC Alain fixed a bug introduced in the activity performed in the mornig which may have caused the unlocks; relock is in progress.

 

TCS

WI and BCK CO2 chillers setpoint change

 

Images attached to this report
AdV-DAQ (Data collection)
poulton - 14:20 Tuesday 21 May 2024 (64313) Print this report
Comment to Glitches in Hrec latency (64233)

Looking more at the CALI latency spikes in the attached plot, I show the CALI, FbmMain and FDS_Img latencies around when the latency spikes happened. The first and the last spikes come from FbmMain and then FDS_Img. There was an issue with FDS_Img when these latency spikes happen when there are issues with FDS_Img. I have also attached the log of FDS_Img when the latency spike occurred.

The second spike is caused by the CALI node when it is in the DOWN state, from the log:

2024-05-11-19h45m38-UTC>INFO...-in DOWN main
2024-05-11-19h45m38-UTC>INFO...-DOWN> enable_actions: 1
2024-05-11-19h45m38-UTC>INFO...-RESET CALnoise process to its default configuration
2024-05-11-19h45m38-UTC>INFO...-cm_send(b'CALnoise', b'AcRelayChTranSet', b'CAL_filter_enbl', -1)
2024-05-11-19h45m38-UTC>INFO...-Waiting 2.0 second
2024-05-11-19h45m40-UTC>INFO...-cm_send(b'CALnoise', b'AcRelayChTranSet', b'PCAL_filter_enbl', -1)
2024-05-11-19h45m40-UTC>INFO...-Memory Used increase, cur. 2259476.00(KB), inc. 6156.00(KB)
2024-05-11-19h45m40-UTC>INFO...-Waiting 2.0 second
2024-05-11-19h45m42-UTC>INFO...-cm_send(b'CALnoise', b'ReloadConfig', b'', b'reset')
2024-05-11-19h45m42-UTC>INFO...-Waiting 1.0 second


The latency spike is caused by CALI's use of the virgotools.wait function in the CALI_Common library rather than self.timer[' XXX']. This causes the latency to spike for 5 seconds every time the CALI node goes into DOWN.

Images attached to this comment
Non-image files attached to this comment
AdV-DAQ (Data collection)
masserot - 12:28 Tuesday 21 May 2024 (64312) Print this report
Comment to Hrec latency : PyDAQ data collection updated (64247)

As trial to reduce the impact of the SQZ_FLT latency fluctuation on the FbmAlp one :

  • the SQZ_FLT server is now sending its frames directly the the SqzMoni server , instead of FbmAlp previously
  • the SQZ_FLT channels are now propagated through the Detector monitoring DAQ branch to be transmited to the FbmSt  frame merger
  • The channel monitoring the SQZ_FLT latency is now the channel called Daq_SqzMoni_SQZ_FLT_latency  (see the plot)

Operation complete at 2024-05-21-10h15m-UTC

Images attached to this comment
AdV-TCS (CO2 laser projector)
taranto - 12:01 Tuesday 21 May 2024 (64311) Print this report
WI and BCK CO2 chillers setpoint change

Since the DMS flag of WI_CO2_Laser became red for mean_TCS_WI_CO2_PWROUT value falling down, I recovered WI CO2 main laser state by acting on its chiller setpoint:

  • at 08.26 UTC WI chiller setpoint changed from 19.015°C to 19.035°C → wrong direction

  • at 08.30 UTC WI chiller setpoint changed from 19.035°C to 19.005°C → right direction

  • at 08.56 UTC WI chiller setpoint changed from 19.005°C to 18.995°C → right direction

The BCK chiller set point has been changed accordingly (18.995°C).

After these actions, the CO2 powers were at their nominal values, without acting on the waveplates.

AdV-TCS (CO2 laser projector)
lorenzini, nardecchia, berni - 11:55 Tuesday 21 May 2024 (64310) Print this report
CO2 piezo-driver test

This morning, we performed some tests on the NI CO2 laser piezo actuator in preparation for the possible CP drum mode (expected at 2.8 kHz (63854)) Q factor measurement.

We injected sine lines at 4 different frequencies (85 Hz, 1040 Hz, 2040 Hz, 3040 Hz) with 4 V amplitude (in fig.1 and fig.2 two examples are shown) measuring the response amplitude on the NI_CO2_ISSIN photodiode. 

We found that the response of the laser power is consistent with a cut-off frequency around roughly 2500 Hz.

At low frequency the measured response is in agreement with the values that we obtained on Feb  13th 2024 (63258). 

Images attached to this report
AdV-DET (Commissioning)
gouaty, masserot, mwas - 11:54 Tuesday 21 May 2024 (64309) Print this report
B1s_PD1_112MHz phi computation

To avoid some jumps on the B1s_PD1_112MHz phase, its computation ha been improved by setting a threshold on its magnitude :

  • SDB2_photodiode.cfg configuration file
  • line 823 commented and replaced by the lines 824 - 827
  • The B1s_PD1_112MHz_mag threshold has been set to 0.0005

The SDB2_photodiodes server has been restarted at 2024-05-21-07h04m43-UTC .  The attached plot shows that with the

  • before the update, the B1s_PD1_112MHz _phi is oscillating between +/-4pi
  • after the update as the B1s_PD1_112MHz_mag is below  the threshold , its phase remains at zero

 

Images attached to this report
Comments to this report:
masserot - 15:10 Tuesday 21 May 2024 (64314) Print this report

A mistake was done for the B1s_PD1_112MHz phase computation , it was put before the rotation with 2xF56MHz LNFS phase instead of after .

This has been fixed at 2024-05-21-12h46m29-UTC by upgrading the SDB2_Photodiodes configuration and the restarting of the server

masserot - 16:20 Thursday 23 May 2024 (64333) Print this report

The B1s_PD1_112MHz_mag threshold has been updated  to 0.005  the 2024-05-21-12h47m16-UTC .

Looking at the B1s_PD1_112MHz_{mag,phi} signal trend , trend plot and its zoom,  one can find

  • a  big jump on the B1s_PD1_112MHz_phi signal  when the B1s_PD1_112MHz mag reached the threshold 
  • but  far enough and short (see the raw plot )before enabling the  phase noise subtraction using the B1s_PD1_112MHz phi channel .

This issue has been understand : it s present only on the ACL_OP_CH keyword using the operator  "mag|phi_umwrap_n", not in the  simple "mag|phi" operator . This issue has been fixed in the Acl release v1r25p17

Images attached to this comment
Search Help
×

Warning

×